OpenStudy (jake124):

1/6 + 3/8 A. 4/6 B. 4/8 C. 4/14 D. 13/24

OpenStudy (mathmale):

Your very first step would be to identify the Lowest Common Denominator (LCD). What is it?

OpenStudy (misssmartiez):

As what mathmale said, you locate the LCD to reconfigure the 6 and 8 to the same values.

OpenStudy (misssmartiez):

On some point in the lines, 6 and 8 can be multiplied and share a common factor. Which one is the LEAST?

OpenStudy (shadowlegendx):

\[\frac{ 1 }{ 6 } = \frac{ numerator }{ denominator }\] \[\frac{ 3 }{ 8 } = \frac{ numerator }{ denominator }\]

OpenStudy (jake124):

6 x 8 = 48

OpenStudy (shadowlegendx):

That is true, but we want the lowest common denominator. There is a number "lower" than 48 that 6 and 8, can both go into.

OpenStudy (shadowlegendx):

The best way to do this, is to take the smallest number and multiply it by increasing numbers from 1 -> and so on 6 x 1 = 6 Can 8 go into 6? no 6 x 2 = 12 Can 8 go into 12? no 6 x 3 = 18 Can 8 go into 18? No 6 x 4 = 24 Can 8 go into 24? You tell me :)

OpenStudy (jake124):

yes

OpenStudy (shadowlegendx):

How many times

OpenStudy (shadowlegendx):

???

OpenStudy (jake124):

8 times

OpenStudy (jake124):

13 @ShadowLegendX

OpenStudy (jake124):

so i think its D

OpenStudy (shadowlegendx):

8 goes into 64, 8 times 8 goes into 24...3 times

OpenStudy (jake124):

so the answer

OpenStudy (shadowlegendx):

Now with this information 6 goes into 24, 4 times 8 goes into 24, 3 times We have our LCD \[\frac{ 1 }{ 6 }(\frac{ 4 }{ 4 }) + \frac{ 3 }{ 8 }(\frac{ 3 }{ 3 })\]

OpenStudy (jake124):

b

OpenStudy (shadowlegendx):

We get, \[\frac{ 4 }{ 24 } + \frac{ 9 }{ 24 } = \frac{ 13 }{ 24 }\]

OpenStudy (jake124):

thanks !!!!

OpenStudy (shadowlegendx):

No problem :)